Determine where the global extrema of  f(x)=3x2/3−2x

on the interval [−1,1] occur.


1
Expert's answer
2022-04-04T17:35:42-0400
"y=3x^{2/3}-2x, \qquad [-1,1]\\
\begin{aligned}
y'(x) = &2x^{-1/3}-2\\
=& \frac{2}{\sqrt[3]{x}}-2
\end{aligned}"

Since "y'(x)" exists for all "x={0}" , the critical numbers of occur when "y'(x)=0" and "x=0"


"\frac{2}{\sqrt[3]{x}}-2=0 \rightarrow \frac{2}{\sqrt[3]{x}}=2\rightarrow \sqrt[3]{x}=1 \rightarrow x=1\\
y(1)=1\\
y(0)=0"

The values at the endpoints of the interval are:


"y(-1)=5"

Comparing these three numbers, we see that the absolute maximum value is "y(-1)=5" and the absolute minimum value is "y(0)=0"